Amon Amarth Co. offers 10-year, 8 percent coupon bonds with semiannual payments and a yield to maturity of 8.24 percent. What is the market price of a $1,000 face value bond?

Group of answer choices

A. $591.04

B. $1,108.16

C. $990.32

D. $1,521.75

E. $983.86
Answer #1

Number of periods = 10 * 2 = 20

Rate = 8.24% / 2 = 4.12%

Coupon = (0.08 * 1000) / 2 = 40

Market price = Coupon *[1- 1 / (1 + r)n] / r + FV/ (1 + r)n

Market price = 40 * [1- 1 / (1 + 0.0412)20] / 0.0412 + 1000 / (1 + 0.0412)20

Market price = 40 * [1 - 0.44598] / 0.0412 + 445.98147

Market price = 40 * 13.44705 + 198.89947

Market price = $983.86